# Backreaction from quantum fluxes at the Kerr inner horizon

### v1

@@ V Discussion
To understand how the Kerr geometry will react to the diverging quantum field near the inner horizon, it is natural to work in the framework of the semiclassical Einstein field Eqs. (1), wherein the spacetime’s curvature is sourced by the vacuum expectation value of the renormalized stress-energy tensor (RSET) of the quantum field. The potential problems (mathematical, physical, and philosophical) associated with such an approach have been debated again and again over the years [43, 44, 45, 46], but it nonetheless remains true that the semiclassical approach is perfectly valid as an effective field theory of quantum gravity, as long as the RSET remains below Planck-scale energies (at $10^{19}$ GeV, which is already orders of magnitude above the grand unified scale, beyond the point where every atomic nucleus has been dissociated and the very notion of particles and interactions is called into question).
Here it is found that the RSET in double-null coordinates contains a non-zero component ${\langle U|\hat{T}_{vv}|U\rangle_{\text{ren}}}$ near the Kerr inner horizon, which, when converted to coordinates that are regular across that horizon, yields an exponential divergence in the quantum flux. Such behavior was also found in Refs. [14, 17]; the additional contributions here were to compute more inner-horizon RSET components over the full parameter space ($a$, $\theta$) (Fig. 5), find the asymptotic behavior at both high and low black hole spins (Figs. 3, 4), and analyze the backreaction that the RSET elicits.
@@
In semiclassical backreaction analyses, the above statement is usually the end of the story. Ref. [17] takes it one step farther to show that in the inevitable regions where the local geometry neither contracts nor expands, the $v\varphi$-component of the RSET will cause an initial tendency for the geometry to undergo an infinite twisting. However, the question of whether anything more can be ascertained from the RSET about the geometry’s evolution and the black hole’s final configuration is either left to speculation or completely ignored, and for good reason—as soon as the inner horizon is pushed even slightly away from its Kerr value, the Kerr vacuum RSET is no longer valid and a new RSET in a more general spacetime would need to be found to continue the evolution any further. Nonetheless, the goal of the current work is to provide a path forward to do exactly that, so that one may determine whether or not the semiclassical inner-horizon instability generically leads to a spacelike singularity.
@@
The backreaction analysis of Sec. IV.3 employed the conformally separable framework of Refs. [4, 5, 6, 7] to determine how semiclassical fluxes can be understood together with the classical fluxes that lead to mass inflation near the inner horizon. One can immediately gather from the presence of a non-vanishing RSET component ${\langle T^{-}_{rt}\rangle}={\left(\langle T^{-}_{vv}\rangle-\langle T^{-}_{uu}\rangle\right)R^{-2}\Delta_{r}^{-1}}$ that the semiclassical backreaction drives the evolution away from many of the potential (meta)stable geometries proposed as potential endpoints of black hole evolution, like the regular “eye of the storm” geometry or the Kerr “black-bounce” geometry (see the end of Sec. IV.2). In contrast, the conformally separable model generically predicts the presence of a strong, spacelike singularity that is stable over astrophysical time scales.
@@
In particular, in Sec. IV.3, the metric is generalized to allow for all the functions in the Kerr line element of Eq. (II) to respond freely to the presence of semiclassical stress-energy. The dominant evolution near the inner horizon comes from the horizon function $\Delta_{r}(r)$ and the conformal factor $\rho(r,\theta,t)$, and thus it becomes useful to analyze the spacetime in a tetrad frame rotated to align with the principal null directions so that the $\Delta_{r}$- and $\rho$-dependent portions of the vierbein are purely diagonal, Eqs. (16a)–(16d).
@@
In this tetrad frame, the RSET’s dominant components are the diagonals ${\langle T^{-}_{00}\rangle}$ and ${\langle T^{-}_{11}\rangle}$, Eqs. (17), while the off-diagonal ${\langle T^{-}_{01}\rangle}$ also diverges as $\Delta_{r}^{-1}$ but is numerically seen to be usually several orders of magnitude smaller than the diagonal components. As such, one is justified in treating the initial semiclassical evolution with the inflationary Kasner spacetime [8], which uses precisely the dynamical vierbein of Eqs. (16a)–(16d) sourced by stress-energy components $T_{00}$ and $T_{11}$. Eventually (or even initially, in the latitudes of infinite quantum twisting), this solution will break down once the RSET’s angular components become non-negligible, but even once the inflationary Kasner model breaks down, the full conformally separable model can still be continued (albeit with less symmetry), as it has been shown to be robust to a wide variety of stress-energy sources [5].
@@
The advantage of the conformally separable model at describing the inner-horizon backreaction is not only in its ability to encompass a wide range of behaviors related to the mass inflation phenomenon, but also in its potential to take the semiclassical evolution even deeper. It is almost miraculous that such a chaotic instability leads at least initially to a spacetime (the inflationary Kasner metric) that is so simple, elegant, and possesses a high degree of symmetry. The RSET in this new homogeneous spacetime should be relatively straightforward to compute, with the only difficulty coming from mode-matching to the Kerr Unruh state (though the relevant mathematical details have already been worked out in, e.g., Ref. [47]). In fact, the renormalized vacuum polarization $\langle\hat{\phi}^{2}\rangle_{\text{ren}}$ has already been calculated in the inflationary Kasner spacetime in an adiabatic vacuum state [48], with the conclusion that the semiclassical contribution follows the same qualitative behavior as the classical stress-energy.
@@
It is time to take RSET calculations beyond vacuum black holes. From this work it can be gleaned that the semiclassical inner-horizon instability causes the local Kerr spacetime to be filled with diverging streams of radiation, and while this work provides evidence that those streams should generically lead to an astrophysically stable, chaotic, spacelike singularity, their exact semiclassical evolution remains an open but tractable question.

### v2

@@ V Discussion
To understand how the Kerr geometry will react to the diverging quantum field near the inner horizon, it is natural to work in the framework of the semiclassical Einstein field equations [Eq. (1)], wherein the spacetime’s curvature is sourced by the vacuum expectation value of the renormalized stress-energy tensor (RSET) of the quantum field. The potential problems (mathematical, physical, and philosophical) associated with such an approach have been debated again and again over the years [43, 44, 45, 46], but it nonetheless remains true that the semiclassical approach is perfectly valid as an effective field theory of quantum gravity, as long as the RSET remains below Planck-scale energies (at $10^{19}$ GeV, which is already orders of magnitude above the grand unified scale, beyond the point where every atomic nucleus has been dissociated and the very notion of particles and interactions is called into question).
Here it is found that the RSET in double-null coordinates contains a nonzero component ${\langle U|\hat{T}_{vv}|U\rangle_{\text{ren}}}$ near the Kerr inner horizon, which, when converted to coordinates that are regular across that horizon, yields an exponential divergence in the quantum flux. Such behavior was also found in Refs. [1, 17]; the additional contributions here were to compute more inner-horizon RSET components over the full parameter space ($a$, $\theta$) (Fig. 5), find the asymptotic behavior at both high and low black hole spins (Figs. 3 and 4), and analyze the backreaction that the RSET elicits.
@@
In semiclassical backreaction analyses, the above statement is usually the end of the story. Reference [17] takes it one step further to show that in the inevitable regions where the local geometry neither contracts nor expands, the $v\varphi$-component of the RSET will cause an initial tendency for the geometry to undergo an infinite twisting. However, the question of whether anything more can be ascertained from the RSET about the geometry’s evolution and the black hole’s final configuration is either left to speculation or completely ignored, and for good reason—as soon as the inner horizon is pushed even slightly away from its Kerr value, the Kerr vacuum RSET is no longer valid and a new RSET in a more general spacetime would need to be found to continue the evolution any further. Nonetheless, the goal of the current work is to provide a path forward to do exactly that, so that one may determine whether or not the semiclassical inner-horizon instability generically leads to a spacelike singularity.
@@
The backreaction analysis of Sec. IV.3 employed the conformally separable framework of Refs. [5, 6, 7, 8] to determine how semiclassical fluxes can be understood together with the classical fluxes that lead to mass inflation near the inner horizon. One can immediately gather from the presence of a nonvanishing RSET component ${\langle T^{-}_{rt}\rangle}={\left(\langle T^{-}_{vv}\rangle-\langle T^{-}_{uu}\rangle\right)R^{-2}\Delta_{r}^{-1}}$ that the semiclassical backreaction drives the evolution away from many of the potential (meta)stable geometries proposed as potential end points of black hole evolution, like the regular “eye of the storm” geometry or the Kerr “black-bounce” geometry (see the end of Sec. IV.2). In contrast, the conformally separable model generically predicts the presence of a strong, spacelike singularity that is stable over astrophysical timescales.
@@
In particular, in Sec. IV.3, the metric is generalized to allow for all the functional degrees of freedom in the Kerr line element of Eq. (II) to respond freely to the presence of semiclassical stress energy. The dominant evolution near the inner horizon comes from the horizon function $\Delta_{r}(r)$ and the conformal factor $\rho(r,\theta,t)$, and thus it becomes useful to analyze the spacetime in a tetrad frame rotated to align with the principal null directions so that the $\Delta_{r}$- and $\rho$-dependent portions of the vierbein are purely diagonal, Eqs. (16a)–(16d).
@@
In this tetrad frame, the RSET’s dominant components are the diagonals ${\langle T^{-}_{00}\rangle}$ and ${\langle T^{-}_{11}\rangle}$ [Eq. (17)], while the off-diagonal ${\langle T^{-}_{01}\rangle}$ also diverges as $\Delta_{r}^{-1}$ but is numerically seen to be usually several orders of magnitude smaller than the diagonal components. As such, one is justified in treating the initial semiclassical evolution with the inflationary Kasner metric [9], which uses precisely the dynamical vierbein of Eqs. (16a)–(16d) sourced by stress-energy components $T_{00}$ and $T_{11}$. Eventually (or even initially, in the latitudes of infinite quantum twisting), this solution will break down once the RSET’s angular components become non-negligible, but even once the inflationary Kasner model breaks down, the full conformally separable model can still be continued (albeit with less symmetry), as it has been shown to be robust to a wide variety of stress-energy sources [6].
@@
The advantage of the conformally separable model at describing the inner-horizon backreaction is not only in its ability to encompass a wide range of behaviors related to the mass inflation phenomenon, but also in its potential to take the semiclassical evolution even deeper. It is almost miraculous that such a chaotic instability leads at least initially to a spacetime (the inflationary Kasner metric) that is so simple, elegant, and possesses such a high degree of symmetry. The RSET in this new homogeneous spacetime should be relatively straightforward to compute, with the only difficulty coming from mode-matching to the Kerr Unruh state (though the relevant mathematical details have already been worked out in Ref. [47]). In fact, the renormalized vacuum polarization $\langle\hat{\phi}^{2}\rangle_{\text{ren}}$ has already been calculated in the inflationary Kasner spacetime in an adiabatic vacuum state [48], with the conclusion that the semiclassical contribution follows the same qualitative behavior as the classical stress energy.
@@
It is time to take RSET calculations beyond vacuum black holes. From this work, it can be gleaned that the semiclassical inner-horizon instability causes the local Kerr spacetime to be filled with diverging streams of radiation, and while this work provides evidence that those streams should generically lead to an astrophysically stable, chaotic, spacelike singularity, their exact semiclassical evolution remains an open but tractable question.